# Bilayer Excitons in the Laughlin Fractional Quantum Hall State

### v2

@@ II Supplementary Text
II.1 Fractional charge and statistics of excitons belonging to ${}^{2}_{0}$ family
@@
where $\sigma_{i}=\pm 1$. The matrices ${\bf I}$ and ${\bf J}$ are identity and a constant matrix of ones respectively. Since $\dim{\bf K}_{i}$ represent the number of $U(1)$ gauge fields in the low energy Chern-Simons description, we can set $\dim{\bf K}_{i}=|\nu_{i}^{*}|$. Note that this identification only works for the Abelian states when $\nu_{i}^{*}\in\mathbb{Z}$ which we focus on in this section. Similarly, $\sigma_{i}$ is chosen depending upon the ${}^{2}_{0}$ particle-/hole-doped $\Lambda$ levels so we identify $\sigma_{i}=\text{sign}(\nu_{i}^{*})$.
@@
The minimal solutions $\ell_{i}$ can also be used to compute the exciton statistics. Before we compute the exciton statistics, we notice that due to the definition $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$, there is an ambiguity in defining the excitation corresponding to the integer vector ${\bf l}_{i}$. This ambiguity suggests that there are distinct solutions of ${\bf l}$ that all correspond to the same exciton charge, or for that matter, a generic quasiparticle charge. However, to make sure all these distinct solutions describe the same particle, we need to check if all such excitations result in the same mutual and self statistical phases modulo $2\pi$. We proved that the distinct solutions indeed result in the same statistical phases in our previous report (see Ref. Zhang2025fractionalexciton ) for a general bilayer ${}^{a}_{b}$ family. This guarantees that for any choice of ${\bf l}_{i}$ corresponding to the solutions obtained $\ell_{i}$ indeed describes the same minimal exciton as with any other choice of ${\bf l}_{i}$ preserving $\ell_{i}$. We thus set $\left[{\bf l}_{1}\right]_{1}=\ell_{1}$ and $\left[{\bf l}_{2}\right]_{1}=\ell_{2}$, and all the remaining elements are set to zero $[{\bf l}_{i}]_{n}=0$ $\forall~{}n\in[2,|\nu_{i}^{*}|]$. The exciton self-statistics is computed as,
@@
which shows that the excitons, in general, follow anyonic statistics rather than the usually anticipated bosonic character. Interestingly, we notice that since $\nu_{i}^{*}\in\mathbb{Z}$, the sum $2x\equiv\sum_{i=1}^{2}\left(|2\nu_{i}^{*}+1|-2\right)\left(2\nu_{i}^{*}+1\right)\in 2\mathbb{Z}$ for some integer $x\in\mathbb{Z}$. Similarly, the denominator $12y+1\equiv\text{gcd}(|2\nu_{1}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}\in 2\mathbb{Z}+$ 1 for some $y\in\mathbb{Z}$. Now since the following equation has no integer solutions,
@@
Figure S1: Activated behavior of quasiparticle excitations. (a) Arrhenius plot of bulk conductance measured in the parallel flow $G_{\text{PF}}$ (black) and counterflow $G_{\text{counter}}$ (blue) configurations at $\nu_{1}=\nu_{2}=1/3$ and $B$ = 28 T. The activated behavior of $G_{\text{PF}}$ is directly linked to the charge gap, $\Delta_{\text{charge}}$, which represents the energy required to create thermally excited quasiparticles and quasiholes. Interestingly, the activated behavior of $G_{\text{counter}}$ reveals the energy required to generate a charge-neutral excitation, $\Delta_{\text{exciton}}$, which closely aligns with $\Delta_{\text{charge}}$. This provides direct evidence that both quasiparticles and the neutral mode are generated through the same thermal process.

### v3

@@ II Supplementary Text
II.1 I. Fractional charge and statistics of excitons belonging to $^{2}_{0}$CF family
@@
where $\sigma_{i}=\pm 1$. The matrices $\bf I$} and $\bf J$} are identity and a constant matrix of ones respectively. Since $\dim{\bf K}_{i}$ represents the number of $U(1)$ gauge fields in the low energy Chern-Simons description, we can set $\dim{\bf K}_{i}=|\nu_{i}^{*}|$. Note that this identification only works for the Abelian states when $\nu_{i}^{*}\in\mathbb{Z}$ which we focus on in this section. Similarly, $\sigma_{i}$ is chosen depending upon the $^{2}_{0}$CF particle-/hole-doped $\Lambda$ levels so we identify $\sigma_{i}=\text{sign}(\nu_{i}^{*})$.
@@
The minimal solutions $\ell_{i}$ can also be used to compute the exciton statistics. Before we compute the exciton statistics, we notice that due to the definition $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$, there is an ambiguity in defining the excitation corresponding to the integer vector ${\bf l}_{i}$. This ambiguity suggests that there are distinct solutions of $\bf l$} that all correspond to the same exciton charge, or for that matter, a generic quasiparticle charge. However, to make sure all these distinct solutions describe the same particle, we need to check if all such excitations result in the same mutual and self statistical phases modulo $2\pi$. We proved that the distinct solutions indeed result in the same statistical phases in our previous report (see Ref. Zhang2025fractionalexciton) for a general bilayer $^{a}_{b}$CF family. This guarantees that for any choice of ${\bf l}_{i}$ corresponding to the solutions obtained $\ell_{i}$ indeed describes the same minimal exciton as with any other choice of ${\bf l}_{i}$ preserving $\ell_{i}$. We thus set $\left[{\bf l}_{1}\right]_{1}=\ell_{1}$ and $\left[{\bf l}_{2}\right]_{1}=\ell_{2}$, and all the remaining elements are set to zero ${\bf l}_{i}]_{n}=0$ $\forall~n\in[2,|\nu_{i}^{*}|]$. The exciton self-statistics is computed as,
@@
which shows that the excitons, in general, follow anyonic statistics rather than the usually anticipated bosonic character. Interestingly, we note that since $\nu_{i}^{*}\in\mathbb{Z}$, the sum $2x\equiv\sum_{i=1}^{2}\left(|2\nu_{i}^{*}+1|-2\right)\left(2\nu_{i}^{*}+1\right)\in 2\mathbb{Z}$ for some integer $x\in\mathbb{Z}$. Similarly, the denominator $2y+1\equiv\text{gcd}(|2\nu_{1}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}\in 2\mathbb{Z}+1$ for some $y\in\mathbb{Z}$. Now, since the following equation has no integer solutions,
@@
II.2 II. Drag current in the generalized (111) condensate phase
We follow the previous analysis (see Supplementary Information in Zhang2025fractionalexciton) to compute the drag current arising solely from the condensate phase. A few clarifications are in order. By ‘drag current due to the condensate phase’, we mean the contribution generated by the backflow of the condensed ground state. This backflow can be induced by interlayer flux attachments, which induce electric fields in the drag layer. In this situation, a perfect drag response (i.e., $j_{\text{drag}}=-j_{\text{drive}}$) is not guaranteed and will depend on the ground state, and our goal in this section is to quantify precisely this deviation. In contrast, a distinct mechanism, arising from quasiparticle/quasihole pairs forming an exciton between the two layers, yields a perfect drag response whenever excitons move as a whole and do not separate into independent opposite charges. This latter mechanism is not the focus of this section.
Our starting point is the $\bf K$} matrix given in Eq. (M7) in Methods. One way to interpret the $\bf K$} matrix is to treat the four $\Lambda$ levels as separate layers, each with different Hall response and conductance. Our goal is to compute the current density in drag layer $j_{\text{drag}}$ as a function of current density in drive layer $j_{\text{drive}}$. We apply an external electric field along $x$ direction in the drive layer, that should be the same in two $\Lambda$ levels corresponding to the drive layer. The perpendicular external magnetic field is assumed to be pointing in the $z$ direction, and therefore the attached fluxes point along negative $z$ direction.
Since the induced electric field due to the motion of fluxes is ${\bf E}_{\text{ind}}=-{\bf v}\times{\bf B}$, the induced electric field in the $i$th layer is, $\delta E_{ix}=-\sigma_{0}^{-1}p_{ij}j_{jy}$ where $\sigma_{0}=e^{2}/h$ is the conductance quantum. The interlayer and intralayer fluxes are denoted by $p_{ij}$. Similarly, the $y$ component of the induced electric field is given by, $\delta E_{iy}={\color[rgb]{0,0,0}\sigma_{0}^{-1}}p_{ij}j_{jx}$. For the state described by $\bf K$}-matrix Eq. (M7) the flux attachments are $p_{ii}=2$ for all layers, and the interlayer fluxes are $p_{12}=p_{34}=2$ and $p_{23}=1$.
We now compute the current density in each layer due to the external applied electric field and the contributions form the induced electric field. We keep the longitudinal conductance $\sigma^{i}_{xx}\neq 0$ and assume rotational symmetry, thus $\sigma^{i}_{xx}=\sigma^{i}_{yy}$, and $\sigma^{i}_{xy}={\color[rgb]{0,0,0}-\sigma^{i}_{yx}}$, for all $i=1,\dots,4$, we get,
$j_{1x}$ $=\sigma^{1}_{xx}\left[E-2\sigma_{0}^{-1}\left(j_{1y}+j_{2y}\right)\right]-2\sigma_{xy}^{1}\sigma_{0}^{-1}\left(j_{1x}+j_{2x}\right),$ (S10a)
$j_{2x}$ $=\sigma^{2}_{xx}\left[E-\sigma_{0}^{-1}\left(2j_{1y}+2j_{2y}+j_{3y}\right)\right]-\sigma_{xy}^{2}\sigma_{0}^{-1}\left(2j_{1x}+2j_{2x}+j_{3x}\right),$ (S10b)
$j_{3x}$ $=-\sigma^{3}_{xx}\sigma_{0}^{-1}\left(j_{2y}+2j_{3y}+2j_{4y}\right)-\sigma_{xy}^{3}\sigma_{0}^{-1}\left(j_{2x}+2j_{3x}+2j_{4x}\right),$ (S10c)
$j_{4x}$ $=-2\sigma^{4}_{xx}\sigma_{0}^{-1}\left(j_{3y}+j_{4y}\right)-2\sigma_{xy}^{4}\sigma_{0}^{-1}\left(j_{3x}+j_{4x}\right),$ (S10d)
$j_{1y}$ $=2\sigma_{xx}^{1}\sigma_{0}^{-1}\left(j_{1x}+j_{2x}\right)+\sigma_{xy}^{1}\left[E-2\sigma_{0}^{-1}\left(j_{1y}+j_{2y}\right)\right],$ (S10e)
$j_{2y}$ $=\sigma_{xx}^{2}\sigma_{0}^{-1}\left(2j_{1x}+2j_{2x}+j_{3x}\right)+\sigma_{xy}^{2}\left[E-\sigma_{0}^{-1}\left(2j_{1y}+2j_{2y}+j_{3y}\right)\right],$ (S10f)
$j_{3y}$ $=\sigma_{xx}^{3}\sigma_{0}^{-1}\left(j_{2x}+2j_{3x}+2j_{4x}\right)-\sigma_{xy}^{3}\sigma_{0}^{-1}\left(j_{2y}+2j_{3y}+2j_{4y}\right),$ (S10g)
$j_{4y}$ $=2\sigma_{xx}^{4}\sigma_{0}^{-1}\left(j_{3x}+j_{4x}\right)-2\sigma_{xy}^{4}\sigma_{0}^{-1}\left(j_{3y}+j_{4y}\right).$ (S10h)
The drive and drag currents are defined as $j_{\text{drive}}=j_{1x}+j_{2x}$ and $j_{\text{drag}}=j_{3x}+j_{4x}$. The above expressions can be simplified by substituting the quantized value of the Hall conductance of each layer as $\sigma^{i}_{xy}\sigma_{0}^{-1}\equiv\nu_{i}$, where $\nu_{i}=1$ for all $i$. Next, from Ohm’s law $j_{x}\sim\sigma_{xx}E$, therefore, we can rewrite these coupled equations up to first order in $\alpha_{i}$, where we define $\alpha_{i}\equiv\sigma_{xx}^{i}\sigma_{0}^{-1}$.
$j_{1x}$ $=\alpha_{1}\left(E\sigma_{0}-2j_{1y}-2j_{2y}\right)-2\left(j_{1x}+j_{2x}\right),$ (S11a)
$j_{2x}$ $=\alpha_{2}\left(E\sigma_{0}-2j_{1y}-2j_{2y}-j_{3y}\right)-\left(2j_{1x}+2j_{2x}+j_{3x}\right),$ (S11b)
$j_{3x}$ $=-\alpha_{3}\left(j_{2y}+2j_{3y}+2j_{4y}\right)-\left(j_{2x}+2j_{3x}+2j_{4x}\right),$ (S11c)
$j_{4x}$ $=-2\alpha_{4}\left(j_{3y}+j_{4y}\right)-2\left(j_{3x}+j_{4x}\right),$ (S11d)
$j_{1y}$ $=\left(E\sigma_{0}-2j_{1y}-2j_{2y}\right)+\mathcal{O}(\alpha_{1}^{2}),$ (S11e)
$j_{2y}$ $=\left(E\sigma_{0}-2j_{1y}-2j_{2y}-j_{3y}\right)+\mathcal{O}(\alpha_{2}^{2}),$ (S11f)
$j_{3y}$ $=-\left(j_{2y}+2j_{3y}+2j_{4y}\right)+\mathcal{O}(\alpha_{3}^{2}),$ (S11g)
$j_{4y}$ $=-2\left(j_{3y}+j_{4y}\right)+\mathcal{O}(\alpha_{4}^{2}).$ (S11h)
The solution for the transverse current density is $j_{1y}=j_{4y}=E\sigma_{0}/8$, $j_{2y}=5E\sigma_{0}/16$, and $j_{3y}=-3E\sigma_{0}/16$. Plugging these into the expressions for longitudinal current, we obtain,
$j_{1x}$ $=\alpha_{1}\frac{E\sigma_{0}}{8}-2\left(j_{1x}+j_{2x}\right),$ (S12a)
$j_{2x}$ $=\alpha_{2}\frac{5E\sigma_{0}}{16}-\left(2j_{1x}+2j_{2x}+j_{3x}\right),$ (S12b)
$j_{3x}$ $=-\alpha_{3}\frac{3E\sigma_{0}}{16}-\left(j_{2x}+2j_{3x}+2j_{4x}\right),$ (S12c)
$j_{4x}$ $=\alpha_{4}\frac{E\sigma_{0}}{8}-2\left(j_{3x}+j_{4x}\right).$ (S12d)
The above equations can now be solved for the longitudinal current. The solution is,
$j_{1x}=$ $\frac{3}{32}\sigma^{1}_{xx}E-\frac{25}{128}\sigma_{xx}^{2}E-\frac{9}{128}\sigma_{xx}^{3}E-\frac{1}{32}\sigma^{4}_{xx}E,$ (S13a)
$j_{2x}=$ $-\frac{5}{64}\sigma_{xx}^{1}E+\frac{75}{256}\sigma^{2}_{xx}E+\frac{27}{256}\sigma_{xx}^{3}E+\frac{3}{64}\sigma_{xx}^{4}E,$ (S13b)
$j_{3x}=$ $\frac{3}{64}\sigma_{xx}^{1}E-\frac{45}{256}\sigma_{xx}^{2}E-\frac{45}{256}\sigma_{xx}^{3}E-\frac{5}{64}\sigma_{xx}^{4}E,$ (S13c)
$j_{4x}=$ $-\frac{1}{32}\sigma_{xx}^{1}E+\frac{15}{128}\sigma_{xx}^{2}E+\frac{15}{128}\sigma_{xx}^{3}E+\frac{3}{32}\sigma_{xx}^{4}E.$ (S13d)
This solution now can be used to find the longitudinal drag/drive currents given by the definition $j_{\text{drive}}\equiv j_{1x}+j_{2x}$ and $j_{\text{drag}}\equiv j_{3x}+j_{4x}$. We have,
$j_{\text{drive}}$ $=\frac{1}{64}\sigma_{xx}^{1}E+\frac{25}{256}\sigma_{xx}^{2}E+\frac{9}{256}\sigma_{xx}^{3}E+\frac{1}{64}\sigma_{xx}^{4}E,$ (S14a)
$j_{\text{drag}}$ $=\frac{1}{64}\sigma_{xx}^{1}E-\frac{15}{256}\sigma_{xx}^{2}E-\frac{15}{256}\sigma_{xx}^{3}E+\frac{1}{64}\sigma_{xx}^{4}E.$ (S14b)
We can now simplify this further by imposing that the conductance of the first and fourth layer is the same $\sigma_{xx}^{1}=\sigma_{xx}^{4}\equiv\sigma^{\text{I}}_{xx}$. Similarly, the second and third layer have the same conductance as well $\sigma_{xx}^{2}=\sigma_{xx}^{3}\equiv\sigma^{\text{II}}_{xx}$. We get,
$j_{\text{drive}}$ $=\frac{4}{128}\sigma_{xx}^{\text{I}}E+\frac{17}{128}\sigma_{xx}^{\text{II}}E,$ (S15a)
$j_{\text{drag}}$ $=\frac{4}{128}\sigma_{xx}^{\text{I}}E-\frac{15}{128}\sigma_{xx}^{\text{II}}E.$ (S15b)
One cannot explicitly assume that the longitudinal conductance of the two $\Lambda$-levels to be the same. The drag ratio is,
$\frac{j_{\text{drag}}}{j_{\text{drive}}}=\frac{4\sigma_{xx}^{\text{I}}-15\sigma_{xx}^{\text{II}}}{4\sigma_{xx}^{\text{I}}+17\sigma_{xx}^{\text{II}}}.$ (S16)
Physically at low temperatures, we expect $\sigma^{\text{I}}_{xx}\ll\sigma^{\text{II}}_{xx}$. However, the state described by Eq. (M7) does not exhibit perfect drag, therefore, one cannot attribute the perfect drag response solely due to the condensate.
Figure S1: Evolution of FQH states across $\nu_{total}-\Delta\nu$ map. Parallel-flow conductance $G_{\mathrm{PF}}$ measured in the Corbino device as functions of total filling factor $\nu_{\mathrm{total}}$ and $\Delta\nu$ at $B=12$ T (left) and $18$ T (right). The schematic diagram in the bottom panels highlight the observed FQH states at each $B$ field. Composite fermions with 2+1 flux attachment dominates at $B=12$ T, whereas the Jain sequence states dominate at $B=18$ T.
Figure S2: Excitonic pairing between single-component FQH states. (a) Schematic phase diagram of a quantum Hall bilayer as a function of $\nu_{\mathrm{total}}$ and $\Delta\nu$. Red and blue lines show the the expected trajectories for the single-component FQH states of the top and bottom layers, respectively. (b) Parallel-flow conductance $G_{\mathrm{PF}}$ (top), counterflow conductance $G_{\mathrm{counter}}$ (middle), and drag ratio $I_{\mathrm{drag}}/I_{\mathrm{drive}}$ (bottom) measured at fixed $\nu_{1}=1/3$ as functions of $\nu_{2}$ for $B=12$ T and $T=20$ mK in the Corbino device. This measurement corresponds to the same sweep shown in Fig. 3c, but acquired at a lower temperature.
Figure S3: Equal density line in the Corbino device. Parallel-flow conductance $G_{\mathrm{PF}}$ (top), counterflow conductance $G_{\mathrm{counter}}$ (middle), and drag ratio $I_{\mathrm{drag}}/I_{\mathrm{drive}}$ (bottom) measured as functions of the total filling factor $\nu_{\mathrm{total}}$ at $B=31$ T along the equal-density line. This dataset is identical to that shown in Fig. 4b–d, but here we adjust the vertical scales of $G_{\mathrm{PF}}$ and $G_{\mathrm{counter}}$ to highlight their distinct behaviors, which together give rise to the observed perfect drag response.
Figure S4: Activated behavior in counterflow conductance. Arrhenius plot of the counterflow conductance from the Corbino device as a function of temperature at $\nu_{\mathrm{total}}=-2/3$, $\Delta\nu=0$, and $B=28$ T, demonstrating activated behavior. We note that both unpaired charges and excitons contribute to the counterflow conductance. As a result, the observed activation behavior does not directly reflect the energy cost associated with generating excitons.